OpenStudy (anonymous):

I am completely lost on this problem. Assume x = x(t) and y = y(t). Find dx/dt if x^2(y-6) = 12y+3 and dy/dt = 2 when x=5 and y=12. Can you take me through this step by step?

terenzreignz (terenzreignz):

It might be prudent to first try to find \[\Large \frac{dx}{dy}\] implicitly or otherwise :)

OpenStudy (anonymous):

where do you start with the dx/dy, then? Is it d/dx (2x) , d/dx (y-6) , and d/dx (12y-3) ??

terenzreignz (terenzreignz):

Oh, you do it implicitly. At worst, you have to do this... \[\Large x^2(y-6 )=12y-3\\\Large x^2=\frac{12y-3}{y-6}\\\\\\\\\\\Large x = \sqrt{\frac{12y-3}{y-6}}\] And differentiate x with respect to y.

terenzreignz (terenzreignz):

But I think implicit is nicer...